Kolmogorov-Smirnov Tests and Inputs

to HWW

Richard St. Denis University of Glasgow/CDF 19 November 2007

H → WW Meeting

[figure]

LRHWW: No NN cut

Richard St. Denis, Glasgow University/CDF, 19 November 1

H → WW Meeting

[figure]

LRHWW: 0 jets

Richard St. Denis, Glasgow University/CDF, 19 November 2

H → WW Meeting

[figure]

LRHWW: 1 jet

Richard St. Denis, Glasgow University/CDF, 19 November 3